Three cards are drawn randomly from a deck of cards one after the other with replacement each time. Find the probability that the three cards are a 5, a heart and a king.

a. 0.0156
b. 0.125
c. 0.0015
a. 0.0192
1
Expert's answer
2020-12-25T14:11:31-0500

We will use the Multiplication Rule (https://www.statisticshowto.com/multiplication-rule-probability/) to solve the problem. We assume that the deck has 52 cards. The probability that a drawn card is 5 is: "p_1=\frac{4}{52}=\frac{1}{13}." The probability that a card is a king is: "p_2=\frac{4}{52}=\frac{1}{13}" and the probability we get a heart: "p_3=\frac{1}{4}". As a result, we get: "p=p_1\cdot p_2\cdot p_3=\frac{1}{13}\cdot\frac{1}{13}\cdot\frac{1}{4}\approx0.0015"

Answer: 0.0015